OpenStudy (anonymous):

solve for x x=log^3 1/81

OpenStudy (anonymous):

Do you mean \[x=\log_3 (81)\] ?

OpenStudy (anonymous):

yes except it is 1/81 I couldn't get the little three

OpenStudy (anonymous):

Okay. Well, \[\log_3(\frac{1}{81}) = \log_3(\frac{1}{3^4})=\log_3(3^{-4})=-4\]